Anderson International Limited is evaluating a project in Erewhon. The project will create the following cash flows:

Year Cash Flow
0 –$875,000   
1 306,900   
2 287,600   
3 285,300   
4 259,300   

All cash flows will occur in Erewhon and are expressed in dollars. In an attempt to improve its economy, the Erewhonian government has declared that all cash flows created by a foreign company are “blocked” and must be reinvested with the government for one year. The reinvestment rate for these funds is 4 percent. Assume Anderson uses a required return of 10 percent on this project.

a. What is the NPV of the project? (A negative answer should be indicated by a minus sign. Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)
b. What is the IRR of the project? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)
Answer #1

-875000 306900 287600 285300 259300 Anderson International Cash Flows Required return NPV of the project IRR of the project $

Formulae

A =C1565+1 306900 =D1565+1 287600 =E1565+1 285300 =F1565+1 259300 -875000 1565 Anderson International 1566 Cash Flows 1567 Re
Answer #2

NPV= 

(16,393.54)

IRR= 9.38%

answered by: umber
Answer #3

Discount rate 10%   Reinvested with government for 1 year at 4%

CF 0  -875000

CF 1   0

CF 2   306900*1.04=319,176

CF 3   287600*1.04=299,104

CF 4   285300*1.04= 296,712

CF5    259300*1.04=269,672

NPV=(16,393.54)

IRR= 9.38%

answered by: anonymous